A large circus ride has a centripetal acceleration at 6.0 m/s2. If the circular ride has a radius of 11.0 m, what is its velocity at the edge?
Answer #1

   \(Centripetal acceleration = v^{2}/r =6\)   

=>    \(v^{2} = 6*11 =66\)   

=>     \(v =\sqrt{66} =8.12 m/s\)   
Answer #2

centripetal acceleration = v^2/r=6


v=sqrt(66)